OpenStudy (anonymous):

Use the Distributive Property to multiply. 5(b + 8)

Parth (parthkohli):

\( \color{Black}{\Rightarrow (5 \cdot b) + (5 \cdot 8) }\)

OpenStudy (phi):

here is an example, but using numbers 5(6+8)= 5*6+5*8